Rony walks 9 km due east and the 12 km bue south. How far is she from the starting point?
A)18 b)15 c)20 d)12

Answer by algebrahouse.com(1659)

Rony walks 9 km due east and the 12 km bue south. How far is she from the starting point?

A right triangle is formed with legs of 9 and 12.
The hypotenuse {side across from the right angle} is the distance from the starting point.

Pythagorean Theorem
a² + b² = c² {a and b are the legs, c is the hypotenuse}

9² + 12² = c² {substituted into the pythagorean theorem}
81 + 144 = c² {squared the 9 and the 12}
225 = c² {combined like terms}
c = 15 {took the square root of both sides}

15 km from the starting point
www.algebrahouse.com